OpenStudy (anonymous):

Tan^2×sin^2×=tan^2×-sin^2 ×

OpenStudy (anonymous):

\[Tan^2×Sin^2×=Tan^2×-Sin^2 ×\] \[Tan^2x=\frac{Sin^2x}{Cos^2x}\] substitute, \[Tan^2x \times Sin^2x=Tan^2x-Sin^2x --->\frac{Sin^2x}{Cos^2x}\times Sin^2x=\frac{Sin^2x}{Cos^2x}-Sin^2x\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x}{Cos^2x}-SIn^2x\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x}{Cos^2x}-\frac{Sin^2x}{1}\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x}{Cos^2x}-\frac{Sin^2x \times Cos^2x}{Cos^2x}\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x-Sin^2x \times Cos^2x}{Cos^2x}\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x(1-Cos^2x)}{Cos^2x}\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^2x(Sin^2x)}{Cos^2x}\] \[\frac{Sin^4x}{Cos^2x}=\frac{Sin^4x}{Cos^2x}\] \[Identity\]